OpenStudy (anonymous):

find an equation of the tangent line to the curve y=3tanx at the point (pi/4,3) Thanks

OpenStudy (lopezking1):

y = 3 tan x differentiating w.r.t x dy/dx = 3 Sec^2 x which is the slope of the curve slope value m = 3 Sec^2 pi/4 = 3 ( sqr rt 2^2) = 3 x 2 = 6 Equation to tangent @ ( pie/4,3) will be y -- y' = m ( x -- x' ) y' = 3 and x' = pie/4 y -- 3 = 6 ( x -- pie/4 ) ( 6 x -- y) = ( 6 pie -- 3 ) and this is the solution to your question

OpenStudy (anonymous):

Thank You very much

OpenStudy (anonymous):

what does "--" indicate... is that minus?